# Exponential Sums with Sparse Polynomials and Distribution of the Power Generator

### v1

@@ 3.1. Discrepancy and exponential sums
One of the basic tools used to study uniformity of distribution is the celebrated Koksma–Szüsz inequality [20, 36] (see also [9, Theorem 1.21]), which generalises the Erdős–Turán inequality to arbitrary dimensions. Like the Erdős–Turán inequality, the Koksma–Szüsz inequality gives a bound on the discrepancy via certain exponential sums.
@@ 3.2. Poisson summation
We formulate the following result, which is well known and with various modifications has been used in a large number of works. However, since finding a precise general reference is not easy, we present it here with a full proof, which follows closely the argument of Friedlander and Iwaniec [12, Section 3].
@@
As in [12, Section 3], consider the function
@@ 3.3. Multiplicative orders modulo divisors
There is no doubt that the following elementary statement is well-known, it is also given in [25, Lemma 6].
@@
The following lower bound follows immediately from Lemma 3.3, and is also given as [25, Lemma 7].
@@
Next, we apply Lemma 3.3 to derive the following variant of [25, Lemma 7].
@@ 3.4. Bounding exponential sums
is bounded by $O\left(N\left(NT/(R\tau)+1\right)\right)$. Hence, using this for all divisors $R\mid T$ with $R>p^{16/23}$ using the classical bound on the divisor function (see [19, Equation (1.81)]), we see that the number $U$ of pairs $(k,m)$, $1\leq k,m\leq K$, with
@@ 4. Multidimensional Generalisations
As in [25, Section 3.1], we consider polynomials
@@
The following generalisation of the power generator has been suggested in [25, Section 4]. Let $F_{1}=X_{1}^{e}L_{1}+L_{0}$, where $L_{0},L_{1}$ and the remaining polynomials $F_{2},\ldots,F_{m}$ are all non-constant linear polynomials in ${\mathbb{F}}_{p}[X_{2},\ldots,X_{m}]$.
@@
which provides the most natural generalisation of the power generator but also inherits its negative aspects such as homogeneity, see [25, Sections 1.2 and 4].
The argument of [25] allows to get an explicit estimate the discrepancy of the corresponding sequences of vectors only for $s=1$, see, for example, [15, Theorem 3.2] or [25, Corollary 9], and appeal to the result of Bourgain [3, Theorem 1] (that is, to (1.6)), otherwise.

### v2

@@ 5.1. Discrepancy and exponential sums
One of the basic tools used to study uniformity of distribution is the celebrated Koksma–Szüsz inequality [22, 40] (see also [11, Theorem 1.21]), which generalises the Erdős–Turán inequality to arbitrary dimensions. Like the Erdős–Turán inequality, the Koksma–Szüsz inequality gives a bound on the discrepancy via certain exponential sums.
@@ 5.2. Poisson summation
We formulate the following result, which is well known and with various modifications has been used in a large number of works. However, since finding a precise general reference is not easy, we present it here with a full proof, which follows closely the argument of Friedlander and Iwaniec [14, Section 3].
@@
As in [14, Section 3], consider the function
@@ 5.3. Multiplicative orders modulo divisors
There is no doubt that the following elementary statement is well-known, it is also given in [27, Lemma 6].
@@
The following lower bound follows immediately from Lemma 5.3, and is also given as [27, Lemma 7].
@@
Next, we apply Lemma 5.3 to derive the following variant of [27, Lemma 7].
@@ 5.4. Bounding exponential sums
is bounded by $O\left(N\left(NT/(R\tau)+1\right)\right)$. Hence, using this for all divisors $R\mid T$ with $R>p^{16/23}$ using the classical bound on the divisor function (see, for example, [21, Equation (1.81)]), we see that the number $U$ of pairs $(k,m)$, $1\leq k,m\leq K$, with
@@ 6. Multidimensional Generalisations
As in [27, Section 3.1], we consider polynomials
@@
The following generalisation of the power generator has been suggested in [27, Section 4]. Let $F_{1}=X_{1}^{e}L_{1}+L_{0}$, where $L_{0},L_{1}$ and the remaining polynomials $F_{2},\ldots,F_{m}$ are all non-constant linear polynomials in ${\mathbb{F}}_{p}[X_{2},\ldots,X_{m}]$.
@@
which provides the most natural generalisation of the power generator but also inherits its negative aspects such as homogeneity, see [27, Sections 1.2 and 4].
The argument of [27] allows to get an explicit estimate the discrepancy of the corresponding sequences of vectors only for $s=1$, see, for example, [17, Theorem 3.2] or [27, Corollary 9], and appeal to the result of Bourgain [3, Theorem 1] (that is, to (1.7)), otherwise.